OpenStudy (anonymous):

Solve and describe the steps you took to solve: -6

OpenStudy (anonymous):

Isolate \(n\) by adding \(9\) \[ 3 < n < 229 \]

OpenStudy (anonymous):

Find the average of the boundaries: \[ \frac{229+3}{2} = \frac{232}{2} = 116 \]

OpenStudy (anonymous):

Subtract the average: \[ -113<n-116<113 \]

OpenStudy (anonymous):

Change to absolute value: \[ |n-116|<113 \]

OpenStudy (anonymous):

Okay, thank you.